OpenStudy (anonymous):

find all real numbers that satisfy the equation. 2sinx= sqrt3

OpenStudy (shinalcantara):

\[2 sinx = \sqrt{3}\] \[sinx = \frac{ \sqrt{3} }{ 2 }\] get the angle using the unit circle